In 2002 a cable company charged $34.99 per month. In 2005 it cost $43.43.
What is the company's equation? Use C for the cost per month and t for time in years since 2002. (2003=1)

Found 2 solutions by josmiceli, stanbon:
Answer by josmiceli(19441)
Answer by stanbon(75887)

In 2002 a cable company charged $34.99 per month. In 2005 it cost $43.43.
What is the company's equation? Use C for the cost per month and t for time in years since 2002. (2003=1)
----
slope = 12(43.43-34.99)/(4-1) = 12*8.44/3 = 33.76 (amt. of cost gain per year)
-----
intercept:: C(1 yr.) = 12*34.99 so C(0) = 12*34.99-33.76 = 386.12
-----
Equation::
C(t) = 33.76t + 386.12
-------------
Cheers,
Stan H.
--------------------